OpenStudy (anonymous):

Find dy/dx of (1/4)(2^(x))

OpenStudy (gorv):

\[\frac{ 1 }{ 4 }*2^x=\frac{1 }{ 4 }*\frac{ d 2^x}{ dx}\]

OpenStudy (gorv):

=(2^x *ln2)/4

OpenStudy (anonymous):

If you're not familiar with the formula for the derivative, it can always be derived: \[\begin{align*} \frac{1}{4}2^x&=\frac{2^x}{2^2}\\\\ &=2^{x-2}\\\\ &=e^{\ln2^{x-2}}\\\\ &=e^{(x-2)\ln2}\\\\ \frac{d}{dx}\left[\cdots\right]&=\ln2~e^{(x-2)\ln2}\\\\ &=\ln2~e^{\ln2^{x-2}}\\\\ &=\ln2~2^{x-2}\\\\ &=\frac{\ln2}{4}~2^x\\\\ &=\frac{\ln2}{4}~2^x \end{align*}\]